OpenStudy (anonymous):

a parabola is given by the function f(x)=x^2+4x+3. what is the line of symetry for this parabola?

OpenStudy (kropot72):

Find the value of x that makes f(x) a maximum or a minimum. f'(x) = 2x + 4 f'(x) = 0 when x = -2 The line of symmetry is give by the equation x = -2 which is a vertical line parallelto the y axis

OpenStudy (anonymous):

thanks!

OpenStudy (kropot72):

No problem :)

OpenStudy (anonymous):

There is a calculus free approach for this too. For any parabola of the form ax^2+bx+c =0 the line of symmetry is given by \(y = -\frac b {2a} \)

OpenStudy (anonymous):

Revert if you want to know how to prove this.